OpenStudy (anonymous):

establish the identity 2 csc Ɵ/sec Ɵ + 3cos Ɵ/ sin Ɵ = 5 cot Ɵ

OpenStudy (anonymous):

this is the exact question attached

mathslover (mathslover):

\[\Huge{\frac{2csc \theta}{sec\theta} +3\frac{cos\theta}{sin\theta}=5cot\theta}\] \[\Huge{\frac{2}{sin\theta}/\frac{1}{cos\theta}+3\frac{cos\theta}{sin\theta}=5cot\theta}\]

mathslover (mathslover):

\[\huge{\frac{2cos\theta}{sin\theta}+3\frac{cos\theta}{sin\theta}=5cot\theta}\] \[\huge{5\frac{cos\theta}{sin\theta}=5cot\theta}\] \[\huge{5cot\theta=5cot\theta}\]

mathslover (mathslover):

hence proved

OpenStudy (anonymous):

thanks!!! soo much